OpenStudy (anonymous):

Someone help me please! The following function defines a recursive sequence: f(0) = -5 f(1) = 20 f(n) = -4•f(n -1) - 3•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -5, -20, -65, -200, … -5, 20, -92, 372, … -5, -24, -92, -372, … -5, 20, -65, 200, …

OpenStudy (anonymous):

Let the function be f(x) If x = 2, then f(2) = -4.f(2-1) - 3.f(n-2) => f(2) = -4.(20) - 3.(-5) = -80 + 15 = -65

OpenStudy (anonymous):

Understand how you do it? @kame324

OpenStudy (anonymous):

Kind of.. so it could be either a or d @Brainybeauty

OpenStudy (anonymous):

Yes so can you try to find the next term by the same method? I'm here if you need help Just try to do it yourself first :)

OpenStudy (anonymous):

@kame324